Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
bersanetti, tapia, vardaro - 0:08 Wednesday 09 October 2019 (47157) Print this report
ITF Working point tuning an ISC Noise Injection

We moved the cut-off frequency of the Butterworth filter of B4 from 1500Hz to 4000Hz (virgoData/VirgoOnline/SSFS_Filters.cfg)

## DAQ Filter
ACL_FILTER_SET             "DAQ_FILTER_10KHZ"    1    0     0     10
ACL_FILTER_BUTTERWORTH     "DAQ_FILTER_10KHZ"    "lowpass" 8 3    4000

ACL_FILTER_SET             "ERR_I_FTL_10KHZ"    1    0     0     10
ACL_FILTER_BUTTERWORTH     "ERR_I_FLT_10KHZ"    "lowpass" 8 3    4000

ACL_FILTER_SET             "ERR_Q_FLT_10KHZ"    1    0     0     10
ACL_FILTER_BUTTERWORTH     "ERR_Q_FLT_10KHZ"    "lowpass" 8 3    4000

ITF Alignment

We reach LOW_NOISE_3_SQZ and we tuned the ITF alignment working point:

  • DIFFp_TX = -0.01 (from -0.022)
  • DIFFp_TY = 0.0 (from -0.001)
  • PR_TX = 0.018 (from 0.017)
  • BS_TX = -0.001 (from -0.0015)
  • BS_TY = 0.001 (from 0.0025)

Squeezer Shutter cloed:

We noticed a bump in the sensitivity around 20Hz and we tried to understand if it is caused by scattered light from the squeezer (Fig 1)

  • yellow hrec after the closure of the SQZ shutter
  • blue hrec before the closure of the SQZ shutter
  • magenta hrec with the SQZ shutter closed

between 10 and 20 Hz the sensitivity seems to be slightly spoiled by the scattered light from the squeezer. To be sure the test can be repeated in the future.

 

After the alignment activity and the activity on SQZ shutter we left the ITF for the INJ noise injection.
 

Squeezer Phase Scan:

We tuned the optimal phase of the squeezer in order to measure the squeezing level in Hrec (2650-3140Hz) before the increase of Laser power (fig 2)

Optimal phase 0.6 deg, Slow CC phase set point 0.1e-3V

Squeezing level -2.3 dB Antisqueezing level 4.6 dB

 

Noise Injections LN3

  • MICH GPS: 1254594655 duration 120 s
  • SSFS HIGH Freq: GPS 1254595528, duration 120 s, GAIN 20000
  • SSFS LOW  Freq: GPS 1254596451, duration 120 s, GAIN 4000
  • All longitudinal DOFs started at GPS 1254597074
  • All angular DOFs started at GPS 1254598223
  • BS TX: GPS 125460261, GAIN 300e-6 (before 12e-6)
  • BS TY: GPS 1254603172, GAIN 300e-6 (before 12e-6)

Noise Injections LN1

  • SSFS HIGHF: GPS 1254600475, duration 120s, GAIN 20000
  • SSFS LOW  Freq: GPS 1254600618, duration 120 s, GAIN 4000
Images attached to this report
Comments to this report:
masserot - 8:50 Wednesday 09 October 2019 (47161) Print this report

The cut-off frequency of the Butterworth filter of B4 has been moved from 1500 to 4000Hz  in the SSFS_filters.cfg file only but not yet applied at the SSFS_Ctrl server level .

Here the report of the commannd "cm send -to SSFS_Ctrl -type AcParamGet -t  ERR_Q_FLT_10KHZ -handler AcParam" performed at Wed Oct  9 08:45:58 CEST 2019 LT

Message type AcParam received from SSFS_Ctrl
  -text ERR_Q_FLT_10KHZ
  -text
# name  - rampTime(s) - gain - @Freq - max Order - mode # current max order - fmnSOS
ACL_FILTER_SET "ERR_Q_FLT_10KHZ" 1 0 0 10 0 # 10 5

# name - fltType, order, attdB@fCut, fCut, nNotch/fBand
ACL_FILTER_BUTTERWORTH "ERR_Q_FLT_10KHZ" "lowpass" 8 3  1500 0

Filter S parameters - a0 a1 a2 / b0 b1 b2
        SOS(00/04) - S space: "ZERO_AdV"@-1Hz Q:0 - "POLE_2ND"@1500Hz Q:2.56292  - Gain: 1 - Coef 1 0 0 / 1 4.13995e-05 1.12579e-08
        SOS(01/04) - S space: "ZERO_AdV"@-1Hz Q:0 - "POLE_2ND"@1500Hz Q:0.899976  - Gain: 1 - Coef 1 0 0 / 1 0.000117896 1.12579e-08
        SOS(02/04) - S space: "ZERO_AdV"@-1Hz Q:0 - "POLE_2ND"@1500Hz Q:0.601345  - Gain: 1 - Coef 1 0 0 / 1 0.000176443 1.12579e-08
        SOS(03/04) - S space: "ZERO_AdV"@-1Hz Q:0 - "POLE_2ND"@1500Hz Q:0.509796  - Gain: 1 - Coef 1 0 0 / 1 0.000208129 1.12579e-08

 

To ease the filters tuning, the SSFS_Ctrl server configuration has been updated to send at 500KHz the channels SSFS_Err_post and SSFS_Err_pre with the _NS suffix.

To take in account these  modifications the SSFS_Ctrl server must be restaretd

 

Search Help
×

Warning

×